OpenStudy (anonymous):

A spinner is divided into 8 equal segments. The segments are labeled with the letters A, B, E, I, L, O, T, U. Match each compound event with its probability. Round to the nearest percent, if necessary. please anyone help math this up for me

OpenStudy (anonymous):

Probability of L and vowel A) twenty-three percent Probability of a letter before J and a letter after R B) two percent Probability of B and B C) thirteen percent Probability of vowel and consonant D) eight percent

OpenStudy (anonymous):

@No.name @SolomonZelman @midhun.madhu1987 @waterineyes

OpenStudy (kirbykirby):

Do you assume that the 1st letter stated actually occurs first? Like for the 1st one, is it: 1st spin: L 2nd spin: vowel or does the order not matter?

OpenStudy (anonymous):

A spinner is divided into 5 equal segments. The segments are labeled with the letters A, E, M, O, T. Match each compound event with its probability. Round to the nearest percent, if necessary.

OpenStudy (anonymous):

Term Definition Probability of vowel and consonant A) twenty-four percent Probability of M and vowel B) twelve percent Probability of a letter before J and a letter after R C) four percent Probability of A and A D) eight percent

OpenStudy (kirbykirby):

Ok actually I calculated the probabilities, I think the order it is written in is what counts. So you have 5 vowels and 3 consonants... so 8 letters. 1) prob of L = 1/8, prob of vowel = 5/8, so 1/8*5/8 = 5/64=0.078 = 7.8% ~ 8% 2) prob of letter before J can be either: A, B, E, I, so the probability is 4/8 prob. of letter after R can be either: T, U, so the probability is 2/8 overall probability is 4/8*2/8 = 1/8=0.125=12.5% ~ 13% 3) probability of B is 1/8, so a second time it's again 1/8, so the probabilty overall is 1/8*1/8 = 1/64 = 0.0156 = 1.56% ~ 2% 4) probability of vowel = 5/8, prob. of consonant = 3/8 so overall 5/8* 3/8 = 15/64 = 0.234 = 23.4% ~ 23%

OpenStudy (anonymous):

what about the second one i put

OpenStudy (kirbykirby):

It's the same idea as the first question

OpenStudy (kirbykirby):

It's the same idea.. if you don't understand something I can walk you through it.

OpenStudy (anonymous):

i understand i just need the answer because i need to finish this in less then 3 min